OpenStudy (anonymous):

derivative of (2x(x)^1/2 +x)/(x^1/2)

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{2x\cdot x^{1/2}+x}{x^{1/2}}\right)'\]Is that the problem?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Hmm it would probably be good to simplify before taking the derivative..

zepdrix (zepdrix):

If you factor an x^(1/2) out of each term in the numerator,\[\Large\bf\sf\left[\frac{x^{1/2}(2x+x^{1/2})}{x^{1/2}}\right]'\]Then you have a nice simplification from there,\[\Large\bf\sf\left[\frac{\cancel{x^{1/2}}(2x+x^{1/2})}{\cancel{x^{1/2}}}\right]'\]And you're left to differentiate this:\[\Large\bf\sf (2x+x^{1/2})'\]

zepdrix (zepdrix):

Lemme know if you're confused on that first step.

OpenStudy (anonymous):

i got it thanks

zepdrix (zepdrix):

ok good.

OpenStudy (anonymous):

yea i got one more problem . can u help me with that ?

zepdrix (zepdrix):

sure

OpenStudy (anonymous):

At a given time the height of a person is 5 feet and it increases at a rate of 1 inch per year. The person's weight is 100 lb and it increases at a rate of 5 lb/year. (a) What is the body fattness (BMI) of the person? (b) Is the person getting slimmer? (c) Assuming that the instantaneous rate of change of BMI does not change over a year, predict the person's BMI one year from now. Recall that the BMI index is the fraction of the body weight measured in kilograms divided by the square of the body height measure in meters.

OpenStudy (anonymous):

can you help me with this ? i didnt get this one at all

zepdrix (zepdrix):

Let's try to interpret the stuff before part a. Height is changing over time. So height is a function of time. As time changes, the height changes accordingly. We'll write a function for the height, our units of t will be years.\[\Large\bf\sf h(t)\]The starting point is 5 feet (that value is not changing, is not directly associated with t in any way). And then the person grows 1 inch (1/12 ft) per year (per t). So the height is the 5 feet + 1 inch per t.

zepdrix (zepdrix):

\[\Large\bf\sf h(t)=5+\frac{1}{12}t\]

zepdrix (zepdrix):

So after 12 years, the person will have grown to 6 feet.

zepdrix (zepdrix):

Does that make sense? The way I modeled the data?

OpenStudy (anonymous):

yea i get the function u come up with h(t)=5+1/12t

zepdrix (zepdrix):

Ok good. Using the rest of the data, do you see how we can write a function for the person's weight? \(\Large\bf\sf w(t)=?\)

OpenStudy (anonymous):

is it 100+ 5/12 sorry for late reply

OpenStudy (anonymous):

100+5/12(t)

zepdrix (zepdrix):

why 5/12? :)

zepdrix (zepdrix):

Remember, for height we were converting inches to ft. So 1in = (1/12)ft. But here they gave us 100 in pounds, and the 5 is also in pounds, yes?

OpenStudy (anonymous):

oh ok so it is 100 +5(t) ?

zepdrix (zepdrix):

Ok good so we've got our functions for height and weight. \[\Large\bf\sf h(t)=5+\frac{1}{12}t\]\[\Large\bf\sf w(t)=100+5t\]

OpenStudy (anonymous):

yea but im still not getting how we calculate body fatnes.

zepdrix (zepdrix):

In part A, they tell us the relationship for BMI.\[\Large\bf\sf BMI\quad=\quad \frac{w(t)}{h(t)^2}\]

zepdrix (zepdrix):

But they made it a little trickier. They're changing the units on us.. grrr.

zepdrix (zepdrix):

Our weight is currently measured in lbs. We need to convert that to kilograms.

OpenStudy (anonymous):

oh so we divide wt by 2.2?

zepdrix (zepdrix):

yes.

zepdrix (zepdrix):

Part a is probably the hardest part. We're looking for a general formula for calculating body fatness. So not at any particular t value.

OpenStudy (anonymous):

then how we do it ?

zepdrix (zepdrix):

\[\Large\bf\sf w(t)=(100+5t)\;lbs\cdot \left(\frac{1kg}{2.2lbs}\right)\]Gives us,\[\Large\bf\sf w(t)=(45.5+2.3t)\;kg\]

zepdrix (zepdrix):

I divided each component by 2.2.

OpenStudy (anonymous):

and height should be in meters

zepdrix (zepdrix):

\[\Large\bf\sf h(t)=\left(5+.083t\right)\;ft\cdot \left(\frac{?m}{?ft}\right)\]What's the conversion for ft to meters? I can't remember..

OpenStudy (anonymous):

1m=3.28 ft

OpenStudy (anonymous):

so ht is 1.52+.0253(t)

zepdrix (zepdrix):

\[\Large\bf\sf h(t)=(1.5+0.025t)\;m\]Mmm ok good!

OpenStudy (anonymous):

ok so for a part what value of t we put or we take derivative of both

OpenStudy (anonymous):

and then punt in that w/h^2 formula

zepdrix (zepdrix):

Ya, I guess the BMI would just be calculated by...\[\Large\bf\sf BMI(t)\quad=\quad \frac{(45.5+2.3t)\;kg}{\left[(1.5+0.025t)\;m\right]^2}\]

OpenStudy (anonymous):

ok ill try

zepdrix (zepdrix):

hmm

zepdrix (zepdrix):

Oh you have to input it into an online thing? Are you allowed multiple guesses?

OpenStudy (anonymous):

i dont think so . and for part a i guess we just divide 45.5 by 1.5^2

OpenStudy (anonymous):

no i have to write and solve everything

zepdrix (zepdrix):

Oh ok, they want us to start at time t=0? Ok ok fair enough :)

OpenStudy (anonymous):

i saw an example and thats what they have done

zepdrix (zepdrix):

What value did you get for part a?

OpenStudy (anonymous):

but how we figure if persin is slimmer or not ? they got 22.11 as their weight was 50 and we 20.22 as our weight was 45.5 and height being same for both

zepdrix (zepdrix):

For part b? So as t increases, is the BMI output getting larger or smaller? It might make sense to just solve part c and then answer part b afterwards.

OpenStudy (anonymous):

yea ill try

OpenStudy (anonymous):

so part c i got 20.33 and part a it is 20.22 and when i calculate the rate of bmi it is 0.348 i dont what i just said , does this make any sense ?